Easy as it might seem, I have forgotten how to take a simple number and turn it into a decimal.

so if I want to turn 392 to the third decimal place what is the easiest way to do this?
Thanks :o)

Answer by mananth(16949)

divide by 1000
392/1000 = 0.392
...
25/100 = 0.25
..
2/10 = 0.2